It takes Jo 3 hours to paint the house and Sam takes 5 hours. How long do they take if they paint together?
Answer by Jstrasner(112)

Hello,
So for this problem we need to change the way we show each rate of work for each painter.
The first painter can paint a house in 3 hours, so he can paint 1/3 of a house per hour.
The second painter can paint a house in 5 hours, so he can paint 1/5 of a house per hour.
We need to add their rates together:
1/3 + 1/5
To do this, we need to find a common denominator (3 x 5 = 15)
So:
5/15 + 3/15 = 8/15 of the job per hour.
So now we need to find how long it will take to do one whole job so we divide the number of hours (15) by the amount of the job done (8):
15/8 = 1.875 hours => 1.9 hours
I hope this helps!